40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 22 23:53:16 2019, gautam, Update, ALS, IR ALS locking attempt ALS_X_outOfLoopnoise.pdfALS_X_inLoopnoise.pdfCARM_loopShape.pdf
    Reply  Wed Jul 24 00:22:24 2019, gautam, Update, ALS, PSL frequency locked to XARM length using ALS ALS_X_noise_POX.pdfBR_comparison.pdfALS_CARM_OLG.pdf
       Reply  Thu Jul 25 12:25:56 2019, gautam, Update, ALS, IR ALSX noise ALS_broadPeaks.pdf
Message ID: 14802     Entry time: Wed Jul 24 00:22:24 2019     In reply to: 14800     Reply to this: 14811
Author: gautam 
Type: Update 
Category: ALS 
Subject: PSL frequency locked to XARM length using ALS 

Summary:

I succeeded in locking the PSL frequency to the XARM cavity length, with 9 pm RMS (Attachment #1) motion below 1 kHz, by actuating on MC2 to change the IMC length. The locks were pretty stable (~20 minutes) - the dominant cause of lockloss was the infamous ETMX drifting problem.

Details:

  1. I did not need to do anything to fix the anomalosly high BR mode coupling I reported yesterday surprise.
    • To test where this could be coming from - I looked at the ALS spectrum again with the XARM length locked to the PSL frequency using POX.
    • Then I compared the spectra with the BR filter in the XARM servo enabled/disabled, see Attachment #2
    • There bounce/roll peak heights even with the BR filter disabled is ~x100 smaller than what I reported yesterday (it remained the case today, because without enabling the BR filter in the CARM servo bank, the TRX level was fluctuating wildly at ~16 Hz). 
  2. The CARM loop (which is what the PSL frequency was slaved to) had ~150 Hz UGF with ~40 degrees phase margin, see Attachment #3.
  3. The quoted RMS sensing noise is if we trust the old POX calibration - may be off by a factor of a few, but probably not an order of magnitude. I'll recalibrate using the free-swinging Michelson technique in the coming days.
  4. The two broad humps in Attachment #1, centered at ~180 Hz and ~300 Hz, are present in the XARM lock as well - so it is somehow imprinted on the arm cavity length. Fixing that will improve the RMS noise performance significantly.

My main motivation here is to make some measurements and investigate the SoCal idea using a toy system, i.e. a single arm cavity controlled using ALS, so that's what Craig and I will attempt next.

Attachment 1: ALS_X_noise_POX.pdf  31 kB  Uploaded Wed Jul 24 01:31:32 2019  | Hide | Hide all | Show all
ALS_X_noise_POX.pdf
Attachment 2: BR_comparison.pdf  75 kB  Uploaded Wed Jul 24 01:31:47 2019  | Show | Hide all | Show all
Attachment 3: ALS_CARM_OLG.pdf  11 kB  Uploaded Wed Jul 24 01:31:55 2019  | Show | Hide all | Show all
ELOG V3.1.3-